# A QUBO Framework for Team Formation

### v1

@@ Abstract
Keywords:
Team Formation Quadratic Binary Optimization (QUBO) Graph Neural Network (GNN) Combinatorial Optimization.
@@ 1 Introduction
Related Work. Our QUBO-based formulation for the TeamFormation problem applies to all variants requiring a balance between coverage and cost. In this way, our work generalizes a lot of existing work on team formation, relates to work on balancing submodular objectives with other objective functions, and extends ideas from QUBO and deep learning. We briefly discuss these connections here, and refer the reader to Appendix 0.A for more details.
We generalize several prior formulations of team formation [16, 18, 25, 34, 35] by incorporating task coverage and a flexible cost definition into a single objective. Furthermore, our GNN-based method is distinct from the deep learning methods [8, 13, 28] used in prior work. The $Cov()$ function in our TeamFormation objective is nonnegative monotone submodular, and depending on the definition of $Cost()$ used, variants of our general problem relate to balancing submodular and other functions [10, 14, 15, 17]. However, our solution framework is general and it does not rely on our objective functions having these properties.
We borrow ideas from prior work to formulate TeamFormation problems as combinatorial optimization using QUBO [11, 19], and use the unbalanced penalization technique [22] to make our formulation more efficient. We also extend ideas from the deep learning combinatorial optimization literature [30, 31] to design our GNN architecture to solve the TeamFormation problem.
@@ 2.1 Team Formation
Task Coverage. Given an assignment $\mathbf{x}$, we define the coverage of task $J$, denoted by $Cov(J\mid\mathbf{x})$, as the number of skills required by $J$ that are covered by the experts assigned to $J$. That is, $Cov(J\mid\mathbf{x})=|(\cup_{i\in\mathbf{x}}X_{i})\cap J|$, with $0\leq Cov(J\mid\mathbf{x})\leq|J|$. We denote the size of $\mathbf{x}_{i}$, i.e., the assignment for task $J_{i}$, by $z_{i}=||\mathbf{x}_{i}||_{1}$. This corresponds to the sum of 1-entries in $\mathbf{x}_{i}$.
@@ 2.2 Quadratic Unconstrained Binary Optimization
We provide all mathematical details to use unbalanced penalization to formulate team formation LPs into QUBO in Section 3 and Appendix 0.B.
@@ 3.1 The TeamFormation Problem
Throughout this section we use the vector $\mathbf{y}=\mathbf{s}\;||\;\mathbf{x}$ which represents the solution to our problems. We call $\mathbf{y}$ the solution vector. This vector is of size $(m+n)$ and is the concatenation of $\mathbf{s}$ and $\mathbf{x}$, where $\mathbf{s}$ is a binary vector that encodes whether a skill $i$ is covered (resp. not covered) by $\mathbf{s}$ when $s_{i}=1$ (resp. $s_{i}=0$). We also use the $(n\times m)$ skill-membership matrix $E$ such that $E(i,j)=1$ (resp. 00) if expert $i$ has (resp. not) skill $j$.
We omit several mathematical details and refer the reader to Appendix 0.B for all the derivations of the QUBO formulations.
@@ 3.2 Max-k-Cover
QUBO Formulation. Let $\mathbf{y}=\mathbf{s}\;||\;\mathbf{x}$, be the $(m+n)$-size solution vector we described above. Now let $\mathbf{c}$ be another $(m+n)$ vector such that $c_{i}=\lambda$ if $i\leq m$ and skill $i\in J$, and $c_{i}=0$ otherwise. Then, the Max-k-Cover problem can be expressed by the following linear program:
@@
In Appendix 0.B.1 we show in detail how to derive penalty matrices $P_{k}$ and $P_{C}$ corresponding to the LP constraints. Then the $(m+n)\times(m+n)$ square matrix $Q=-\text{diag}({\mathbf{c}})-P_{k}+P_{C}$ provides a QUBO formulation of Max-k-Cover, where minimizing ${\mathbf{y}}^{T}Q{\mathbf{y}}$ corresponds to maximizing $F(\mathbf{x})=\lambda Cov(J\mid\mathbf{x})-Cost_{k}(\mathbf{x})$.
@@ 3.3 Coverage-Linear-Cost
QUBO Formulation. Let $\mathbf{y}=\mathbf{s}\;||\;\mathbf{x}$, be the $(m+n)$-size solution vector we described above. Now let $\mathbf{c}$ be another $(m+n)$ vector such that $c_{i}=\lambda$ if $i\leq m$ and skill $i\in J$, $c_{i}=-\kappa_{i-m}$ if $i>m$; recall that $\kappa_{i}$ is the cost of hiring expert $i$ (see Sec. 2). Then Coverage-Linear-Cost can be expressed as:
@@
In Appendix 0.B.2 we show how to create penalty matrices $P_{1}$ and $P_{2}$ to capture the constraints in the LP. Then, the $(m+n)\times(m+n)$ square matrix $Q=-\text{diag}({\mathbf{c}})-P_{1}+P_{2}$ has the property that minimizing ${\mathbf{y}}^{T}Q{\mathbf{y}}$ corresponds to maximizing $F(\mathbf{x})=\lambda Cov(J\mid\mathbf{x})-Cost_{L}(\mathbf{x})$.
@@ 3.4 Coverage-Graph-Cost
QUBO Formulation. We consider the following constrained linear program that encodes the Coverage-Graph-Cost problem:
@@
For the QUBO formulation we need the solution vector $\mathbf{y}$, we defined above. We also need the $(m+n)$ vector ${\mathbf{c}}=(c_{1},\ldots,c_{(m+n)})$, such that $c_{i}=\lambda$ if $i\leq m$ and skill $i\in J$, and $c_{i}=0$ otherwise. Then we compute the $(n\times n)$ matrix $D$ of pairwise distances such that $D(i,j)=d(X_{i},X_{j})$ and add it to the lower-right $(n\times n)$ submatrix of $\text{diag}({\mathbf{c}})$ to obtain $\hat{D}=\text{diag}({\mathbf{c}})+\begin{bmatrix}\mathbf{0}_{m\times m}&\mathbf{0}_{m\times n}\\ \mathbf{0}_{n\times m}&\;D_{n\times n}\end{bmatrix}$ Now, $F(\mathbf{x})=\mathbf{y}^{T}\hat{D}\mathbf{y}$ encodes the Coverage-Graph-Cost objective. In Appendix 0.B.3 we show how to create penalty matrices $P_{1},P_{2}$ to capture the LP constraints, and how the $(m+n)\times(m+n)$ square matrix $Q=-\hat{D}-P_{1}+P_{2}$ provides a complete QUBO formulation of Coverage-Graph-Cost; that is, minimizing ${\mathbf{y}}^{T}Q{\mathbf{y}}$ corresponds to maximizing $F(\mathbf{x})=\lambda Cov(J\mid\mathbf{x})-Cost_{G}(\mathbf{x})$.
@@ 4.2 Graph Neural Networks
We randomly initialize node embeddings for each of the expert and skill nodes, where the dimension of the embeddings is given by the hyperparameter $d_{0}$. We denote the set of $(m+n)$ embeddings by $H^{(0)}=H_{S}^{(0)}\;||\;H_{\mathcal{X}}^{(0)}$, where ||||| | represents concatenation of the $m$ skill embeddings and $n$ expert embeddings.
@@
We call our method QUBO-GNN and visualize the model architecture in Figure 6 in Appendix 0.C. QUBO-GNN is parametrized by several hyperparameters; we include a tabular summary and discussion of these in Appendix 0.C.2.
@@ 5.1 Experimental Setup
Datasets. We evaluate our methods on several real-world datasets also used in past team formation papers: Freelancer, IMDB, Bbsm [1, 24, 25, 35]. We follow the method of [2] and create social graphs with expert coordination costs for our datasets. We provide summary statistics of the datasets in Table 1. Appendix 0.D.1 has detailed descriptions and pre-processing steps of each dataset.
@@
Implementation Details. We used single-process implementations on a 14-core 2.4 GHz Intel Xeon E5-2680 processor for all our experiments. We implement our QUBO-GNN architecture in Python using PyTorch [26] and Deep Graph Library[36], and fine-tune model hyperparameters using grid search. For each dataset, we train separate QUBO-GNN models for up to 100 different tasks. For the normalizing coefficient we set $\lambda=50$, which yields a reasonable balance between weighting coverage and cost for our TeamFormation variants. To aid reproducibility, we report the full set of model parameters used in Appendix 0.D.2, and make our code  [^1] https://github.com/kvombatkere/Team-Formation-QUBO available online.
@@ 5.2 Quantitative Comparison
We evaluate our algorithms against the baselines with respect to our overall objective (and the corresponding coverage, size and cost). We show detailed results for Coverage-Linear-Cost here, and provide full experimental results for Max-k-Cover and Coverage-Graph-Cost in Appendix 0.D. Note that the general experimental patterns observed were similar across all three TeamFormation variants.
@@ 5.3 Transfer Learning
Figure 5 shows a scatter plot of sorted objectives $F$ of QUBO-GNN-Sim and the two baselines for 100 new tasks across each dataset for Coverage-Linear-Cost. The results for the other two problems are shown in Appendix 0.D.5.
@@ 6 Conclusions and Future Work
Future Work. Finding optimal penalty parameters heuristically for our QUBO formulations is challenging, consequently opening up an avenue for future work on efficient methods to tune these penalties. A natural extension of our work would be to consider multiple input tasks and explore more (complex) expert cost functions based on workload, team diameter, etc. Finally, there is scope for fine-tuning the QUBO-GNN model architecture to improve performance both during training and transfer learning.
Appendix
Appendix 0.A Related Work
We present a discussion of prior work in team formation, submodular maximization, QUBO combinatorial optimization and deep learning.
Submodular Maximization. The coverage function is monotone and submodular, which is useful within discrete objective functions, as it encodes a natural diminishing returns property and also comes with an extensive literature on optimization techniques [9, 10, 17]. The greedy algorithm achieves a $1-1/e$ approximation for maximizing a nonnegative monotone submodular function subject to a cardinality constraint [23]. There is also work involving maximizing submodular minus modular or linear functions, where no multiplicative approximation guarantees are possible in polynomial time due to potential negativity [14, 15].
Appendix 0.B QUBO Formulations for TeamFormation
@@ 0.B.1 Max-k-Cover
0.B.2 Coverage-Linear-Cost
@@
0.B.3 Coverage-Graph-Cost
@@
Appendix 0.C QUBO-GNN Model Specifications
0.C.1 QUBO-GNN Model Architecture
Figure 6 illustrates our GNN model architecture. We adopt a two-layer (heterogeneous) graph convolution network (GCN) architecture, with forward propagation given by $H^{(1)}=\sigma_{1}\left(\sum_{r\in\mathcal{R}}\Theta_{r}^{0}H^{(0)}\right)$ and $H^{(2)}=\sigma_{2}\left(\sum_{r\in\mathcal{R}}\Theta_{r}^{1}H^{(1)}\right)$, where $\mathcal{R}$ is the set of different edge types. $H^{(0)}$ represents the input node embeddings of size $d_{0}$, and $H^{(1)}$ and $H^{(2)}$ are the hidden and output layer representations of sizes $d_{h}$ and $(m+n)$, respectively. $\Theta_{r}^{0}$ and $\Theta_{r}^{1}$ are trainable weight matrices specific to $r$, allowing QUBO-GNN to learn different transformations per edge type; $\sigma_{1},\sigma_{2}$ are non-linear activation functions, applied element-wise; we use ReLU for $\sigma_{1}$ and a sigmoid for $\sigma_{2}$.
We include a tabular summary and discussion of the QUBO-GNN model hyperparameters in Section 0.C.2.
[…]
0.C.2 QUBO-GNN Model Hyperparameters
Table 3 provides a summary of the hyperparameters of the QUBO-GNN model, and heuristic ranges of values to grid-search.
Table 3: Description of QUBO-GNN model parameters.
Parameter Description Heuristic range
$p_{1}$ QUBO penalty 1 $[10^{-1},10^{2}]$
$p_{2}$ QUBO penalty 2 $[10^{-1},10^{2}]$
$\lambda$ Normalizing coefficient $[1,10^{2}]$
$d_{0}$ Size of node embeddings $[(m+n)^{1/2},(m+n)/2]$
$d_{h}$ Size of hidden layer $[(m+n)^{1/2},(m+n)/2]$
$p_{d}$ Dropout probability $[0.1,0.3]$
$\alpha$ Binary regularization weight $[1,10]$
$\beta$ Learning rate $[10^{-4},10^{-2}]$
Penalty Multipliers. Capturing problem constraints effectively in a QUBO formulation requires the selection of suitable scalar penalties $p_{1},p_{2}$. In practice, we observed for our problems that the unbalanced penalization scheme yields good solutions for a wide range of values of $p_{1},p_{2}$. However, to enable convergence to better near-optimal solutions we implement a grid search for $p_{1},p_{2}$ over the range of heuristic values shown in Table 3.
Appendix 0.D Additional Experimental Details
0.D.1 Descriptions of Experimental Datasets
@@
0.D.2 QUBO-GNN Implementation Parameters
Table 4 specifies the model parameters used to formulate the three TeamFormation variants into QUBO and instantiate QUBO-GNN to find solutions. Note that for all experiments we used $\lambda=50$. For the QUBO penalty parameters $p_{1},p_{2}$, we grid-searched over the heuristic range $[10^{-1},10^{2}]$ for each training task, and selected the values $p_{1},p_{2}$ that yielded the best task-specific objectives.
@@ 0.D.2 QUBO-GNN Implementation Parameters
0.D.3 Experimental Results for Max-k-Cover
We evaluate our algorithms against the baselines with respect to our overall objective. We observe that Qsolver has the best performance for all datasets for Max-k-Cover. We use $\overline{Cov}=\frac{1}{t}\sum_{i=1}^{t}Cov(J_{i}|\mathbf{x})$ to denote the mean coverage, and $\overline{z}=\frac{1}{t}\sum_{i=1}^{t}z_{i}$ to denote the mean solution size, across training tasks $J_{1},\ldots,J_{t}$. We observe from Table 5 that all three methods find solutions yielding high coverages but Qsolver and Greedy both find the highest values of $\overline{Cov}$. We note that Qsolver often find assignments with a larger solution size than Greedy, but the size of these assignments almost always satisfy the $k$-cardinality constraint; we observed a few violations for Bbsm-3. QUBO-GNN also satisfies the $k$-cardinality constraint but performs between 5-20% worse than Qsolver. This gap in performance could be attributed to insufficient hyperparameter tuning, or suboptimal constraint penalization, since the Qsolver solution indicates that the QUBO formulation is able to find an assignment of experts that balances coverage with the constraints for Max-k-Cover.

### v2

@@ 1 Introduction
2 Related Work
Our QUBO-based formulation for the TeamFormation problem applies to all variants requiring a balance between coverage and cost. In this way, our work generalizes a lot of existing work on team formation, relates to work on balancing submodular objectives with other objective functions, and extends ideas from QUBO combinatorial optimization and deep learning.
Submodular Maximization. The coverage function is monotone and submodular, which is useful within discrete objective functions, as it encodes a natural diminishing returns property and also comes with an extensive literature on optimization techniques [9, 10, 17]. The greedy algorithm achieves a 11/e1-1/e approximation for maximizing a nonnegative monotone submodular function subject to a cardinality constraint [23]. There is also work involving maximizing submodular minus modular or linear functions, where no multiplicative approximation guarantees are possible in polynomial time due to potential negativity [14, 15].
@@ 3.1 Team Formation
[…]
@@ 3.2 Quadratic Unconstrained Binary Optimization
We provide all mathematical details to use unbalanced penalization to formulate team formation LPs into QUBO in Section 4.
@@ 4.1 The TeamFormation Problem
Throughout this section we use the vector $\mathbf{y}=\mathbf{s}\;||\;\mathbf{x}$ which represents the solution to our problems. We call $\mathbf{y}$ the solution vector. This vector is of size $(m+n)$ and is the concatenation of $\mathbf{s}$ and $\mathbf{x}$, where $\mathbf{s}$ is a binary vector that encodes whether a skill $i$ is covered (resp. not covered) by $\mathbf{s}$ when $s_{i}=1$ (resp. $s_{i}=0$). We also use the $(n\times m)$ skill-membership matrix $E$ such that $E(i,j)=1$ (resp. 0) if expert $i$ has (resp. not) skill $j$.
Due to space constraints, we omit several mathematical details and refer the reader to the supplementary material for derivations of the QUBO formulations.
@@ 4.2 Max-k-Cover
QUBO Formulation Sketch. Let $\mathbf{y}=\mathbf{s}\;||\;\mathbf{x}$, be the $(m+n)$-size solution vector we described above. Now let $\mathbf{c}$ be another $(m+n)$ vector such that $c_{i}=\lambda$ if $i\leq m$ and skill $i\in J$, and $c_{i}=0$ otherwise. Then, the Max-k-Cover problem can be expressed by the following linear program:
@@
We derive penalty matrices $P_{k}$ and $P_{C}$ corresponding to the LP constraints. Then the $(m+n)\times(m+n)$ square matrix $Q=-\text{diag}({\mathbf{c}})-P_{k}+P_{C}$ provides a QUBO formulation of Max-k-Cover, where minimizing ${\mathbf{y}}^{T}Q{\mathbf{y}}$ corresponds to maximizing $F(\mathbf{x})=\lambda Cov(J\mid\mathbf{x})-Cost_{k}(\mathbf{x})$.
@@ 4.3 Coverage-Linear-Cost
QUBO Formulation Sketch. Let $\mathbf{y}=\mathbf{s}\;||\;\mathbf{x}$, be the $(m+n)$-size solution vector we described above. Now let $\mathbf{c}$ be another $(m+n)$ vector such that $c_{i}=\lambda$ if $i\leq m$ and skill $i\in J$, $c_{i}=-\kappa_{i-m}$ if $i>m$; recall that $\kappa_{i}$ is the cost of hiring expert $i$ (see Sec. 3). Then Coverage-Linear-Cost can be expressed as:
@@
We create penalty matrices $P_{1}$ and $P_{2}$ to capture the constraints in the LP. Then, the $(m+n)\times(m+n)$ square matrix $Q=-\text{diag}({\mathbf{c}})-P_{1}+P_{2}$ has the property that minimizing ${\mathbf{y}}^{T}Q{\mathbf{y}}$ corresponds to maximizing $F(\mathbf{x})=\lambda Cov(J\mid\mathbf{x})-Cost_{L}(\mathbf{x})$.
@@ 4.4 Coverage-Graph-Cost
QUBO Formulation Sketch. We consider the following constrained linear program that encodes the Coverage-Graph-Cost problem:
@@
For the QUBO formulation we need the solution vector $\mathbf{y}$, we defined above. We also need the $(m+n)$ vector ${\mathbf{c}}=(c_{1},\ldots,c_{(m+n)})$, such that $c_{i}=\lambda$ if $i\leq m$ and skill $i\in J$, and $c_{i}=0$ otherwise. Then we compute the $(n\times n)$ matrix $D$ of pairwise distances such that $D(i,j)=d(X_{i},X_{j})$ and add it to the lower-right $(n\times n)$ submatrix of $\text{diag}({\mathbf{c}})$ to obtain $\hat{D}=\text{diag}({\mathbf{c}})+\begin{bmatrix}\mathbf{0}_{m\times m}&\mathbf{0}_{m\times n}\\ \mathbf{0}_{n\times m}&\;D_{n\times n}\end{bmatrix}$ Now, $F(\mathbf{x})=\mathbf{y}^{T}\hat{D}\mathbf{y}$ encodes the Coverage-Graph-Cost objective. We create penalty matrices $P_{1},P_{2}$ to capture the LP constraints; the $(m+n)\times(m+n)$ square matrix $Q=-\hat{D}-P_{1}+P_{2}$ provides a complete QUBO formulation of Coverage-Graph-Cost; that is, minimizing ${\mathbf{y}}^{T}Q{\mathbf{y}}$ corresponds to maximizing $F(\mathbf{x})=\lambda Cov(J\mid\mathbf{x})-Cost_{G}(\mathbf{x})$.
@@ 5.2 Graph Neural Networks
We randomly initialize node embeddings for each of the expert and skill nodes, where the dimension of the embeddings is given by the hyperparameter $d_{0}$. We denote the set of $(m+n)$ embeddings by $H^{(0)}=H_{S}^{(0)}\;||\;H_{\mathcal{X}}^{(0)}$, where $||$ represents concatenation of the $m$ skill embeddings and $n$ expert embeddings.
@@
[…]
We call our method QUBO-GNN and visualize the model architecture in Figure 2. QUBO-GNN is parametrized by several hyperparameters; Table 1 provides a summary of the hyperparameters of the QUBO-GNN model, and heuristic ranges of values to grid-search.
Table 1: Description of QUBO-GNN model parameters.
Parameter Description Heuristic range
$p_{1}$ QUBO penalty 1 $10^{-1},10^{2}]$
$p_{2}$ QUBO penalty 2 $10^{-1},10^{2}]$
$\lambda$ Normalizing coefficient $1,10^{2}]$
$d_{0}$ Size of node embeddings $(m+n)^{1/2},(m+n)/2]$
$d_{h}$ Size of hidden layer $(m+n)^{1/2},(m+n)/2]$
$p_{d}$ Dropout probability $[0.1,0.3]$
$\alpha$ Binary regularization weight $[1,10]$
$\beta$ Learning rate $10^{-4},10^{-2}]$
Capturing problem constraints effectively in a QUBO formulation requires the selection of suitable scalar penalties $p_{1},p_{2}$. In practice, we observed for our problems that the unbalanced penalization scheme yields good solutions for a wide range of values of $p_{1},p_{2}$. However, to enable convergence to better near-optimal solutions we implement a grid search for $p_{1},p_{2}$ over the range of heuristic values shown in Table 1.
@@ 6.1 Experimental Setup
Datasets. We evaluate our methods on several real-world datasets also used in past team formation papers: Freelancer, IMDB, Bbsm [1, 24, 25, 35]. We follow the method of [2] and create social graphs with expert coordination costs for our datasets. We provide summary statistics of the datasets in Table 2. Detailed descriptions and pre-processing steps of each dataset are available in the supplementary material.
@@
Implementation Details. We used single-process implementations on a 14-core 2.4 GHz Intel Xeon E5-2680 processor for all our experiments. We implement our QUBO-GNN architecture in Python using PyTorch [26] and Deep Graph Library[36], and fine-tune model hyperparameters using grid search. For each dataset, we train separate QUBO-GNN models for up to 100 different tasks. For the normalizing coefficient we set $\lambda=50$, which yields a reasonable balance between weighting coverage and cost for our TeamFormation variants. To aid reproducibility, we report the full set of model parameters used in the supplement, and make our code  [^1] https://github.com/kvombatkere/Team-Formation-QUBO available online.
@@ 6.2 Quantitative Comparison
We evaluate our algorithms against the baselines with respect to our overall objective (and the corresponding coverage, size and cost). Due to space constraints, we only show detailed results for Coverage-Linear-Cost, and provide experimental results for Max-k-Cover and Coverage-Graph-Cost in the supplement. Note that the general experimental patterns observed were similar across all three TeamFormation variants.
@@ 6.3 Transfer Learning
Figure 6 shows a scatter plot of sorted objectives $F$ of QUBO-GNN-Sim and the two baselines for 100 new tasks across each dataset for Coverage-Linear-Cost. The results for the other two problems are shown in the supplement.
@@ 7 Conclusions and Future Work
Future Work. Finding optimal penalty parameters for our QUBO formulations is challenging, consequently opening up an avenue for future work on efficient methods to tune these penalties. A natural extension of our work could consider multiple input tasks and explore more (complex) expert cost functions based on workload, team diameter, etc. Finally, there is scope for fine-tuning the QUBO-GNN model architecture to improve performance.
Acknowledgment. This work was supported by gifts from Microsoft and Google.
Supplementary Material
Appendix 0.A QUBO Formulations for TeamFormation
@@ 0.A.1 Max-k-Cover
0.A.2 Coverage-Linear-Cost
@@
0.A.3 Coverage-Graph-Cost
@@
Appendix 0.B Additional Experimental Details
0.B.1 Descriptions of Experimental Datasets
@@
0.B.2 QUBO-GNN Implementation Parameters
Table 4 specifies the model parameters used to formulate the three TeamFormation variants into QUBO and instantiate QUBO-GNN to find solutions. Note that for all experiments we used $\lambda=50$. For the QUBO penalty parameters $p_{1},p_{2}$, we grid-searched over the heuristic range $10^{-1},10^{2}]$ for each training task, and selected the values $p_{1},p_{2}$ that yielded the best task-specific objectives.
@@ 0.B.2 QUBO-GNN Implementation Parameters
0.B.3 Experimental Results for Max-k-Cover
We evaluate our algorithms against the baselines with respect to our overall objective. We observe that Qsolver has the best performance for all datasets for Max-k-Cover. We use $\overline{Cov}=\frac{1}{t}\sum_{i=1}^{t}Cov(J_{i}|\mathbf{x})$ to denote the mean coverage, and $\overline{z}=\frac{1}{t}\sum_{i=1}^{t}z_{i}$ to denote the mean solution size, across training tasks $J_{1},\ldots,J_{t}$. We observe from Table 5 that all three methods find solutions yielding high coverages but Qsolver and Greedy both find the highest values of $\overline{Cov}$.